OpenStudy (anonymous):

express the following in simplest radical form. square root of 6/27

OpenStudy (anonymous):

\[\sqrt{\frac{ 6 }{ 27 }}=\frac{ \sqrt{6} }{ \sqrt{27} }=\frac{ {\sqrt{2*3}} }{ \sqrt{9*3} }=\frac{ \sqrt{2} }{3 }\]